OpenStudy (anonymous):

When solving x2 – 8x = –11 by completing the square what value is added to both sides of the equation?

OpenStudy (anonymous):

4 but idk if it wil be - or +

OpenStudy (anonymous):

always +

OpenStudy (anonymous):

and it is not 4 you add to both sides, but the square of 4

OpenStudy (anonymous):

\[x^2-8x=-11\]half of 8 is 4 and \(4^2=16\) so next line is \[(x-4)^2=-11+16\]